# Density Ternary Goldbach for primes in a fixed residue class

### v2

@@ 3. Proof of local result
Let $m$ be an odd square free integer divisible by $15$. Let $m=15m^{\prime}$ where $(m^{\prime},30)=1.$ Identifying $\mathbb{Z}_{m}\equiv\mathbb{Z}_{m}^{\prime}\times\mathbb{Z}_{15}$, let $(u,v)\in\mathbb{Z}_{m}$ with $v\equiv 0\pmod{3}$ where $u\in\mathbb{Z}_{m^{\prime}},v\in\mathbb{Z}_{15}$. Define
$f^{\prime}(x)=\frac{2}{\phi(15)}\sum_{y\in\mathbb{Z}_{15}^{*}}f(x,y)$
for $x\in\mathbb{Z}_{m^{\prime}}^{*}$. Observe the sum $\sum_{x\in\mathbb{Z}_{m^{\prime}}^{*}}f^{\prime}(x)>\frac{\phi(m^{\prime})}{2}$ so we may use Proposition 3.1 to conclude that there exists $a_{1},a_{2},a_{3}\in\mathbb{Z}_{m^{\prime}}^{*}$ with $u=a_{1}+a_{2}+a_{3}$ such that
@@
Define $f_{i}:\mathbb{Z}_{15}^{*}\rightarrow[0,1]$ for $i=1,2,3$ by
$f_{i}(y)=f(a_{i},y).$
The condition (3.6) satisfied by the $f^{\prime}(a_{i})$ allows the use of Proposition 3.2 to conclude that there exists $b_{1},b_{2},b_{3}\in\mathbb{Z}_{15}^{*}$ with $v=b_{1}+b_{2}+b_{3},$ such that

### v3

@@ 3. Proof of local result
Let $m$ be an odd square free integer divisible by $3$. Let $m=m_{1}m_{2}$ where $m_{2}$ divides $15$ and $(m_{1},30)=1$. Note that $m_{2}\in\{3,15\}$. Identifying $\mathbb{Z}_{m}\equiv\mathbb{Z}_{m_{1}}\times\mathbb{Z}_{m_{2}}$, let $(u,v)\in\mathbb{Z}_{m}$ where $u\in\mathbb{Z}_{m_{1}},v\in\mathbb{Z}_{m_{2}}$ with $v\equiv 0\pmod{3}$. Define
$f^{\prime}(x)=\frac{2}{\phi(m_{2})}\sum_{y\in\mathbb{Z}_{m_{2}}^{*}}f(x,y)$
for $x\in\mathbb{Z}_{m_{1}}^{*}$. Observe that
$\sum_{x\in\mathbb{Z}_{m_{1}}^{*}}f^{\prime}(x)>\frac{\phi(m_{1})}{2},$
and so we may use Proposition 3.1 to conclude that there exist $a_{1},a_{2},a_{3}\in\mathbb{Z}_{m_{1}}^{*}$ with $u=a_{1}+a_{2}+a_{3}$ such that
@@
Suppose first that $m_{2}=3$. In this case we have $f^{\prime}(a_{i})=f(a_{i},1)$ for $i=1,2,3$. Setting $x_{i}=f(a_{i},1)$ for $i=1,2,3$, Inequality (3.6) becomes
(3.7) $x_{1}x_{2}+x_{2}x_{3}+x_{3}x_{1}>\frac{1}{2}(x_{1}+x_{2}+x_{3}).$
We may equivalently express Inequality 3.7 as
(3.8) $(x_{1}+x_{2}+x_{3})^{2}-(x_{1}+x_{2}+x_{3})>x_{1}^{2}+x_{2}^{2}+x_{3}^{3}.$
Now observe, by an application of Cauchy-Schwarz, we have that
(3.9) $x_{1}^{2}+x_{2}^{2}+x_{3}^{2}\geq\frac{(x_{1}+x_{2}+x_{3})^{2}}{3}.$
Combining Inequalities (3.8) and (3.9) it follows that
$x_{1}+x_{2}+x_{3}>3/2.$
Hence $(u,v)=(u,0)=(a_{1},1)+(a_{2},1)+(a_{3},1)$ with
$f(a_{1},1)+f(a_{2},1)+f(a_{3},1)>3/2.$
Since Inequality 3.7 implies that $x_{i}>0$, it follows that $f(a_{i},1)>0$ for $i=1,2,3$ as desired.
Now suppose $m_{2}=15$. Define $f_{i}:\mathbb{Z}_{15}^{*}\rightarrow[0,1]$ for $i=1,2,3$ by
$f_{i}(y)=f(a_{i},y)$
for each $y\in\mathbb{Z}_{15}^{*}.$ The condition (3.6) satisfied by the $f^{\prime}(a_{i})$ allows the use of Proposition 3.2 to conclude that there exists $b_{1},b_{2},b_{3}\in\mathbb{Z}_{15}^{*}$ with $v=b_{1}+b_{2}+b_{3},$ such that